I'm not really sure how this would be classified...It isn't in my book, it's from a take home study sheet.
Find three consecutive integers such that the sum of the three is 61 more than the largest.....
Thank you very much for any help you can offer.
cory

Answer by mukhopadhyay(490)

Since they are consecutive, assume three numbers as (x-1), x, and (x+1)
Sum of them is 61 more than (x+1)
So, (x-1)+x+(x-1) = 61+(x+1)
=>3x=61+x+1
=>2x=62
=>x=31
The numbers are 30,31, and 32
Verification:
Sum of them = 30+31+32=93
Greatest number+61 = 32+62 = 93 (it matches)
Good luck